STUDIES IN LOGIC, GRAMMAR AND RHETORIC 1 (14)

Halina Świȩczkowska

LANGUAGE AS THE MIRROR OF THE MIND

The Leibnizian attitude toward language is most fully expressed in the conviction that language is the finest mirror of the mind (UG, §1, NE, III, vii, §6) and therefore the analysis of the substance of language should thus lead to a recognition of the mechanisms of the process of thinking itself. Language capacity, in a like manner as the capacity to think, is a function of the same intensity and the same driving force which differentiates Man from other creatures. If this interpretation is on target, it leads us to the conclusion that any natural language system may be the subject of analysis. This results from the fact that every language “reflects” the natural order of ideas which potentially belongs to the intellectual equipment of all substances capable of reflecting about themselves as “Me”. (GP, II, 53; L, 337).

Indeed, this conclusion finds confirmation in many statements made by Leibniz regarding the relation between language and thinking and the knowledge accumulated in the language. The following is a fragment of Analysis Linguarum of 11 September, 1678 in which Leibniz writes:

“Although there are many human languages, all of them sufficiently developed to be suitable for the transmission of any science whatsoever, it is enough, I think, to consider one language: any nation can in fact make discoveries and direct the sciences in its own backyard.” (C, 352; Dascal, 1987, 162).

Consequently, independent of the level of civilizational development of a given language society, it is capable of making discoveries and registering in the language the intellectual process whose aim is to broaden knowledge. Such thesis indeed lies at the basis of classical social anthropology.^{1} Despite the fact that language studied in its historical perspective registers, as Leibniz acknowledges, the history of our discoveries, which is reflected in the example of those “who as Copernicans continue to say that the sun rises and sets” (GP, IV, 459; L, 320); this does not impair Leibniz’s conviction

1 See Malinowski, 1931.

Halina /Swi/eczkowska

that “languages are the best mirror of the human mind, and that a precise analysis of the significations of words would tell us more than anything else about the operations of the understanding”. (NE, III, vii, §6)

The motif of the mirror image appears in a number of Leibniz’s texts. Let us consider how the relationship of reflection is to be understood and what is the role assigned to the “mirror” itself. It may be assumed that signs as well as language are essential to the understanding of the functioning of the mind only when they are the basis of some direct method or when they are the only accessible means of reaching the contents of the mind itself. And also that the mind reflects itself in a mirror, even when this is the only means of learning anything about its nature. This does not lead us to the assumption that this mirror plays a role in the functioning of the mind itself. It seems that Leibniz has gone far from such superficial formulation of the relationship between signs and thinking.^{2} Language in Leibniz is integrally connected with the mind because according to him:

“all human reasoning is performed by means of certain signs or characters. Indeed, it is neither possible nor desirable that the things themselves or even the ideas of them be always distinctly observed by the mind. So, for reasons of economy, signs are used for them”. (GP, VII, 204; Dascal, 1987, 181)

Matters connected with how language involves itself in thinking and what functions are assigned by Leibniz in this process will be considered in a separate study. Below we concentrate on the concept of expression which is fundamental for Leibniz.

Expression

The relation of reflection becomes clear in the context of another, wider concept – “representation” or “expression” – constituting the core of Leibnizian metaphysics. Let us note that: “each created monad represents the whole universe” (GP, VI, 620; L, 649) and “each monad is a living mirror, or a mirror endowed with an internal action, and (...) it represents the universe according to its point of view and is regulated as completely as is the universe itself”. (GP, VI, 598; L, 637) because, since monads are, essentially, closed to any external influence in the process of representing the world, they turn to the content of their interior, finding there the ideas of things.

2 See Dascal, 1987, Foreword, X.

Language as the mirror of the mind

“Expression is common to all the forms and is a genus of which natural perception, animal feeling, and intellectual knowledge are species. In natural perception and feeling it suffices that what is divisible and material and is found dispersed among several beings should be expressed or represented in a single indivisible being or in a substance which is endowed with a true unity. The possibility of such a representation of several things in one cannot be doubted, since our soul provides us with an example of it. But in the reasonable soul this representation is accompanied by consciousness, and it is than that it is called thought.” (GP, II, 112; L, 339)

In Leibniz the notion of expression is a technical concept; therefore, it needs some clarification. In his essay What is an Idea, Leibniz explains that, what is common to different forms of expression is that from the learning of the relations which it expresses we may achieve knowledge of properties corresponding to the thing expressed. The idea is what is in the mind. But as Leibniz writes:

“There are many things in our mind, however, which we know are not ideas, though they would not occur without ideas – for example, thoughts, perceptions, and affections. In my opinion, namely, an idea consists, not in some act, but in the faculty of thinking, and we are said to have an idea of a thing even if we do not think of it, if only, on a given occasion, we can think of it.” (GP, VII, 263-64; L, 207)

According to Nicolas Jolley, Leibniz reduces the discussion of ideas to the discussion of the ways of thinking due to formulation of ideas as disposition to think in a given way.^{3} The idea assumes therefore a certain closer ability or ease of thinking about a thing, the idea of which we have inside, although it not a thought itself. This ability must lead however not only to thinking about the thing, but it has to also express this object. In Discourse on Metaphysics, Leibniz develops this concept by writing:

“As a matter of fact, our soul always does have within it the disposition to represent to itself any nature or form whatever, when an occasion arises for thinking of it. I believe that this disposition of our soul, insofar as it expresses some nature, form, or essence, is properly the idea of the thing, which is in us and is always in us whether we think of it or not.” (L, 320)

The means of expressing things is based on the analogy of relations between a thing and what constitutes its expression. And therefore “speech expresses thoughts and truths, characters express numbers, and an algebraic equation expresses a circle or some other figure”. (GP, VII, 263-64; L, 207) Resemblance is of no need here, for even if the idea of a wheel would not resemble a wheel, it is still possible to elicit from it the truth about the wheel itself.

3 See Jolley, 1990, 162.

The mind therefore thinks about a thing at that moment when its idea is formulated by it in a certain way. A question arises: does such interpretation assume thinking only at the level of consciousness? The idea of a thing summoned by the mind leads however, as Leibniz says, to thinking about the thing itself under one condition: when simultaneously the idea somehow expresses that thing. If an idea is a disposition to think in a particular mode, then these modes of thinking about a thing are connected with a set of possible forms for the expression of a given thing. Indeed the concept of the idea of the thing seems to be in Leibniz integrally connected with a certain skill of expression. If we however are not able to “grasp” the idea of a thing in any form, does this mean that we do not possess this idea? Leibniz would strongly deny this. He would presumably answer that this idea, although present in the mind, is still in a state of “slumber”, is potential, and as such contains all possible forms of expressing things. This results from the fact that

“we have all these forms in our own minds, and even from eternity, for et every moment the mind expresses all its future thought and already thinks confusedly of everything of which it will ever think distinctly.” (L, 320)

It is just a matter of activating an appropriate cognitive process which will release the disposition to reach out for one of the forms of expression. In relation with the above, one can not assume that we are talking about ideas as dispositions to think about things, which is correlated with expressing them: it is always conscious thought. These dispositions may lead to bringing about their actualization and then they may be interpreted as active dispositions. This does not contradict the opinion that the mind of an infant is able to grasp for example the idea of Turing’s universal machine or express the quantum theory of gravitation. In the case of infants as well as the majority of people, these ideas are in the sphere of potentiality, and although it is not known if they will ever achieve disposition to express these ideas on the conscious level, still their minds are programmed to achieve this or any other disposition leading to the expression of idea of things in question. We find the confirmation of such interpretation in the theory of substance. For Leibniz assumes that a given disposition is the result of cooperation of the so- called “minute perceptions”, imperceptible impressions which undergo processing of the mind. (NE, II, i, §15) The disposition, finding a basis in the actions of the mind, leads to expression – the representation of things. Admittedly, the ability for representation is vested in all substances, since “every individual substance in its own manner expresses the universe”. (GP, I, 383-84). But, as we remember, in the thinking soul representation merges with consciousness and only then is it named thinking. (GP, II, 112; L. 339) Such a formulation may lead however to the conclusion that the inner system of representation in any monad is the same system which is vested also in thinking beings – the difference lying only in the fact that the thinking being is able to recognize the system thanks to the ability to think. This conclusion however is false, on the grounds that only minds are vested in the natural system of representation, labeled by Leibniz the natural order of ideas. An indispensable condition of thinking is the existence of something which one may think about. This “something” is for Leibniz the idea itself. Animals, having no ideas, are unable to think. They do however have some system of representation other than ideas, because their souls, just like other monads, express the world.^{4}

Ideas as dispositions

Leibniz’s theory of ideas is interpreted as a disposition theory. Such a manner of interpretation is indicated by numerous statements made by Leibniz, for example the one that ideas and truths are innate in us – as inclinations, dispositions, tendencies or natural potentialities. (NE, Preface, 52). The human mind by its nature contains the basis of certain concepts and theories only to be awakened in appropriate conditions by outside objects. (NE, Preface, 49) Inborn ideas are inborn in the way that the mind draws them from itself; this does not mean that every mind begins as if from the whole stock of ideas and innate truths, or even, that every mind ever reaches the clear knowledge of all those truths which can be drawn from it. According to Leibniz however, there is no such barbarian who, in a matter important to him, would not reflect on the behavior of a liar contradicting himself. (NE, I, i, §4) But such an almost instinctive use of the principle of contradiction does not mean that everyone knows this principle out of necessity, it means only that it is potentially innate, as are other mathematical truths. The proof of learning these truths is to become aware of them and the adequate expression of them. This potential character of innate ideas assumes a parallel invariability of the equipping of the human mind with the invariability of the whole human race. It is also suggested that aboriginal man had at his disposal the same set of truths and ideas

4 Leroy E. Loemker in the introduction to his translation of selected writings of Leibniz (L, 42) expresses the opinion that on the level of unconscious perceptions, Man basically does not differ from other creatures. Thus he accepts that obscure knowledge (cognito obscura) combined with access to unclear notions or ideas is vested also in animals. This however is in contradiction with the Leibnizian assumption concerning the lack of ideas in substances other than the mind.

which also constitute the intellectual stock of people living today. But this conclusion contradicts Leibniz’s principle of individualization. The same equipping of minds results in their lack of differentiation. Leibniz safeguards himself against this consequence by treating ideas as dispositional. This differentiation corresponds to the range of dispositions, and it is possible to treat every single disposition as building from a certain initial state which we may call predisposition that finds its grounding in the unconscious perceptions of the mind and may be reduced to them.^{5} In accordance with this, though the mind of aboriginal man already had the predisposition to discover integral and differential calculus, only in Leibniz do we find that this predisposition has evolved into a full disposition.

Dispositional features of ideas are illustrated by Leibniz by the example of the marble block in which the form of Hercules already exists as if potentially contained within, although the sculptor has to make a big effort to extract it from the block. This allows us to assume that the mind possesses a certain disposition to find in itself such truths and principles for which it is initially programmed, just as the grain of the marble conditions the movement of the sculptor’s chisel. We see here a fairly close analogy with the thesis of Descartes which says that certain ideas are inborn in the same way as certain family diseases, not because the “infants of these families suffer from these diseases in their mother’s womb, but because they are born with a certain disposition or propensity for contracting them”.^{6}

Implicit knowledge

Let us note that in the description of the thesis of disposition, the assumption of Leibniz fulfills a very important role concerning the implicit character of inborn features. The fact that the concepts are inborn implicite in the mind means that “[this] should signify only that the mind has a faculty for knowing them; (...) it has in addition a faculty for finding them in itself, and the disposition, if it is thinking properly, to accept them”. (NE, I, i, §21). Subsequently, the ability of learning notions alone is something different than disposition. As Jolley notices, Leibniz assigns to the concept of implicit knowledge a different sense than to dispositional knowledge. Jolley interprets dispositional knowledge as

5 See Jolley, 1990, 162.

6 R. Descartes, Notes Against a Certain Programme, trans. E. S. Haldane and G. R. T. Ross, in: The Philosophical Works of Descartes, Cambridge University Press, 1911, I, p. 442.

“express knowledge”,^{7} whereas implicit knowledge is knowledge containing, for example, axioms. “The statement that the body is greater than the trunk differs from Euclid’s axiom only in that the axiom restricts itself to precisely what needs to be.” (NE, IV, vii, §10) “Thus, we use these these maxims without having them explicitly in mind. It is rather like the way in which one has potentially in mind the suppressed premisses in enthymemes, which are ommited in our thinking of the argument as well as in our outward expression of it.” (NE, I, i, §4)

Implicit knowledge is therefore, as Leibniz claims, united with the certain ability to learn it or gain it. However, does this ability always lead to its revelation, or does the mind have access to it? We are referring here to Leibniz’s differentiation between truths of reason and truths of fact. Truths of reason are necessary statements in the sense that they are apparent statements or they are reduced to them. Among the truths of reason, Leibniz differentiates primary truths which he names “identities”, “because they seem to do nothing but repeat the same thing without telling us anything”. (NE, IV, ii, §1) Among them he distinguishes two categories: affirmative or negative. Assuming the implicit character of these principles, he describes the mechanism of affirmative statements such as: “an equilateral rectangle is a rectangle” or “a square is not a circle”. Leibniz claims that these statements have features of innate truths because when considering them we implement the principle of identity (and the principle of contradiction respectively), “for in thinking it, one applies the principle of identity and the principle of contradiction to materials which the understanding itself provides.” (NE, I, i, §18). Thus they constitute the exemplification of these innate principles and should not be treated as various truths, instead “regarding the axiom as embodied in the example and as making the example truth” (NE, IV, vii, §10). A child’s acceptance of the statement “the mother is not the father” (meaning that it has the certain disposition to find this truth), consequently reveals the general logical principal of which the above statement is an example. The fact that a child has at its disposal certain implicit knowledge, which consists of, as Leibniz assumes, the principles of classical logic, is not sufficient however to its actualization. Thus it is doubtful if a child, even after long training, could achieve the ability to reveal these principles, though it uses them efficiently on the intuitive level. In other words a child does not possess the disposition to implement them consciously. Despite being devoid of this disposition, it cannot possibly to be able to reveal them.

7 See Jolley, 1984, 172.

By this interpretation, having implicit knowledge is not sufficient to be transformed into disposition connected with its expression. But we should also remember that Leibniz discussing ability, in particular the ability to learn, did not have in mind exclusively “the bare faculty”. As he has written:

“inactive faculties – in short, the pure powers (...) are (...) mere fictions, unknown to nature (...). For where will one ever find in the world a faculty consisting in sheer power without performing any act?” (NE, II, i, §2)

The ability to learn activates itself thanks to the potential characteristic of every mind which Leibniz calls effort (conatus, appetition). Ability joined with this power leads to a certain result – this result being the revelation of an appropriate disposition.

In Meditations on Knowledge, Truth, and Ideas (GP, IV, 422-26, L, 291-94) with reference to the division of ideas in Descartes, Leibniz conducts a detailed classification of ideas from the point of view of the role they play in cognition. The content of this treatise is also repeated in New Essays... in book II, Chapter 29-31. The starting point of this classification is the division of ideas – and the knowledge which corresponds to them – into obscure (cognitio obscura) and clear (cognitio clara). An idea is obscure when it is not sufficient to recognize the thing represented by it. We deal with this then when we consider “some term which the Scholastics had defined poorly, such as Aristotle’s entelechy (...) or other such terms of which we have no sure definition”. (L, 291) We may therefore assume that the mind remains in certain contact with such an idea as the representation of thing; however, this state of knowledge does not lead to the identification of thing. Obscure ideas can be interpreted on the ground of Leibniz’s metaphysics as perceptions which have not yet transformed themselves into aperception, although they may lead to it. Let us point out that Leibniz, accepting the division into obscure and clear ideas as a starting point, assumes at the same time that the mind always finds itself in a certain state of knowledge at which the lowest level marks obscure knowledge and not the lack of it. The mind therefore has to always possess from the beginning of its existence access to the ideas of things which is the consequence of the implicit character of these ideas. In the face of this we may assume that there exist many truths we have access to on the perception level, however the disposition to grasp them is lacking. For “the faculty of knowing innate notions” alone is only a prerequisite to having the “faculty for finding them in itself and the disposition to accept them”.

Disposition and aperception

Leibniz clearly connects the ability of aperception with disposition when he writes that

“to be aware of what is within us, we must be attentive and methodical (...). For thoughts are actions whereas items of knowledge (or truths), in so far as they are within us even when we do not think of them, are tendencies or dispositions; and we know many things which we scarely think about”. (NE, I, i, §25-26)

Undoubtedly, the only means of reaching the content of the mind is to direct it to its natural abilities. Leibniz illustrates this process adducing the Platonic dialogue Menon in which Socrates, guiding a boy by asking questions, helps him to discover in himself innate laws of geometry. (NE, I, i, §5) Aperception alone demands however a certain impetus or a whole series of stimuli which direct the mind to its content – thanks to, for example the properly formulated questions by Socrates. It does not seem probable however that the state of full consciousness of the laws of geometry could be achieved by an infant, and Leibniz did not want to accept “that every innate truth is known always and by everyone”. (NE, I, ii, §11) although cases occur of children reaching these and other laws.^{8} He was protected from such formulation by the assumption concerning the potential and possible features of ideas.

Leibniz in a way confirms the above understanding of the process leading to apercetion by writing that “innate maxims make their appearance only through the attention one gives to them” (NE, I, i, §27) Admittedly every man is equipped with the ability of reaching the innate set of ideas, but the minds of children and savages on which Leibniz bases his examples are less spoiled and corrupted by custom and less molded by the learning which shapes this attention. (NE, I, i, §27) At this point we need to again return to the notion of implicit knowledge. We may assume that this knowledge is vested in every mind. Supposing however it does not become the subject of reflection of the mind – the mind will not activate the disposition to reveal it; this means that it will remain in the area of those activities of the mind which are responsible for thinking on the level of unconsciousness. We are speaking here of activating the disposition, bearing in mind the fact that dispositions to act are based in Leibniz on a whole series of minute perceptions which cause “the mind has a disposition (as much active as passive) to draw (truths) from its own depths”. But for

8 Example supplied by Leibniz in NE, I, i §55.

this to happen certain stimuli are needed “to give the mind the opportunity and the attention for this, and to direct it toward certain necessary truths rather than others”. (NE, I, i, §5) Maybe therefore, implicit knowledge is interrelated with a series of predispositions or passive dispositions which properly stimulated allow the elevation to the state of consciousness. This is what Leibniz has in mind when he writes, “that the less one knows the closer one comes to sharing with blocks of marble and bits of wood the advantage of being infallible and faultless. But unfortunately that is not the respect in which one comes close to them; and so far one is capable of knowledge, it is a sin to neglect to acquire it, and the less instruction one has had the easier it is to fail in this”. (NE, I, i, §27)

Thus Leibniz assumes that although in the initial phase, let us suppose at the time of birth, we are potentially equipped with all possible knowledge and we are ready to obtain it, the learning mechanism activates itself at the moment we receive proper impulses. At the beginning these are impulses originating from the outside – from the senses – because

“children should attend more to the notions of the senses, because attention is governed by need. Hovewer, we shall see later that nature has not ’taken pains to no purpose’ in imprinting us, innately, with items of knowledge; for without these there would be no way of achieving actual knowledge of necessary truths in the demonstrative sciences, or of learning the reasons for facts; and we should have nothing over the bests.” (NE, I, i, §25)

In this context Leibniz’s comments connected with the intellectual capabilities of children become interesting. According to him children in their first years of development do not differ much from animals in the externalized usage of the mind. Therefore, teachers who shape their behavior may implement the same methods as trainers of animals. These tools do not suffice however when we deal with those minds which are capable of taking advantage of them. Taking advantage means achieving the capability of remembering, discovering and judging, and these features are characteristic only of intelligent beings. But since every mind in fact is closed to external influence any impulse originating from the senses according to the principle of pre-establisched harmony corresponds with certain activities of the mind. If we assume that “every action of the mind is thought”, a child’s mind, as opposed to an animal’s, is capable of transferring certain unconscious actions to the conscious level, while at the same time sorting them and finding explanations of them. (AA, VI, i, 275; L, 88) For Leibniz notices that:

“The human mind is analogous to a sieve: the process of thinking consist in shaking it until all the sublest items pass through. Meanwhile as they pass through, Reason acts as an inspector snatching out whatever seems useful.” (C, 170)

What fundamentally differentiates man from animals is the ability to prove, to find rational connections between particular events. I. Hacking writes that “to be human is to be able to prove a little. A man ’who will not reason about anything’ is, as Aristotle says, ’no better than a vegetable’”. (Hacking, 1975, 69) However, the approaching of the content of the mind itself is irrevocably connected, according to Leibniz with the process of the shaping of the mind. He argues this in the following way:

“If someone looks attentively at more pictures of plants and animals than another person, and at more diagrams of machines (...), and if he reads more imaginative novels (...), then he can be said to have more knowledge than the other, even if there is not a word of truth in all that he has seen and heard.” (NE, IV, i, §2)

Knowledge then is a certain process which reveals itself on its particular levels and Leibniz claims with full conviction that for example the knowledge of the rules of classical logic can be achieved only after one acquires the knowledge of mathematics, history and other subjects for how can one order one’s thoughts who has never thought of much? (L, 468) But in another place he also notes that logic is built into thoughts (AA, IV, i, 286; L, 89), and consequently the science of logic is in effect a process of discovering what is already inside of us.

Let us inquire here: on what basis are we able to evaluate how much a child or any other intelligent being knows? A simple answer obviously is: potentially it knows everything. The above question however concerns essentially the dispositional knowledge which is able to activate the given mind. Approaching a possible answer will be the statement that the mind knows that which it is able to grasp in the act of aperception. But the state of gaining knowledge in truth is known only to oneself and the only way to examine this state by another mind is through its externalization; for example, through language. And it seems that essentially this is what Leibniz means when he says that language is the best mirror of the human mind. We may therefore assume that the knowledge which we have at our disposal in an aware or conscious form (active dispositional knowledge) is that part of potential knowledge which is possible to express in some language. Dispositional character of knowledge is closely correlated with the ability of revealing it in the act of speech.

Leibniz clearly indicates that the “activation” of dispositions takes place through the process of education. A Man learning about the world through the senses learns to find a basis for the data referring to the content of his mind. One could obviously assume that everyone of us is capable of discovering many truths independently of any contact with the world through the senses – this is what Leibiz claims with reference to arithmetic and geometry when he writes that

“one could construct these sciences in one’s study and even with one’s eyes closed, without learning from sight or even from touch any of the needed truths; although it is true that if one had never seen or touched anything, one would not bring to mind the relevant ideas. For it is an admirable arrangement on the part of nature that we cannot have abstract thoughts which have no need of something sensible, even if it be merely symbols such as the shapes of letters, or sounds; though there is no necessary connection between such arbitrary symbols and such thoughts”. (NE, I, i, §5)

Focusing on language, Leibniz in his argumentation refers to the metaphysical principle of pre-established harmony, according to which traces of the senses are indispensable for this harmony between the soul and the body to take place. It seems that this argument is of too high a caliber. One could ask how else than through language the process of learning takes place. We must agree here with the view expressed by the contemporary Platonist Hans George Gadamer that consciousness never faces the world without language.

“Wir sind vielmehr in allem Wissen von uns selbst und allem Wissen von der Welt immer schon von der Sprache umgriffen, die unsere eigene ist. Wir wachsen auf, wir lernen die Welt kennen, wir lernen die Menschen kennen und am Ende uns selbst, indem wir sprechen lernen. Sprechen lernen heibt nicht: zur Bezeichnung der uns vertrauten und bekannten Welt in den Gebrauch eines schon vorhandenen Werkzeuges eingefuhrt werden, sondern es heibt, die Vertrautheit und Erkenntnis der Welt selbst, und wie sie uns begegnet, erwerben.” (Gadamer 1967, 96)

This view is in accord with Leibniz’s assumption present in his theory of the mind and his theory of language that linguistic interpretation of the world proceeds any thought and any learning.^{9} Activation of the mind and activation of its disposition are in Leibniz inseparably connected with the existence of language. Language stands at the source of the process of learning which begins at the moment of our birth. It has a very important activating function. Throughout all of our education we are exposed to the impact of the word. Unconsciously immersed in the language we look for things and meanings corresponding to its particular words. At first

“he takes care to put the words in in the order that others customarily do, and contents himself with the thought that he could grasp their sense if the need

9 In this context Leibniz’s comments contained in Préface à la science générale (C, 156) become essential – particularly the one which states that language precedes thought.

arose. Thus a person is sometimes – oftener indeed than he thinks – a mere passer-on of thoughts, a carrier of someone else’s message, as though it were a letter”. (NE, III, ii, §2)

Having however the possibility of “grasping” the thought itself with the assistance of other means of expression than the thought itself, we are capable of reflection on the content of the thought. For as Leibniz writes,

“although characters are arbitrary, their use and connection have something which is not arbitrary, namely a definite analogy between characters and things, and the relations which different characters expressing the same thing have to each other. This analogy or relation is the basis of truth”. (L, 184)

The natural and the psychological order of knowledge

Leibniz makes a clear distinction between the natural and the psychological order of knowledge. He assumes the existence of a natural order of ideas which is “common to angels and men and to intelligences in general”. (NE, III, i, §5) This order being fixed and unchangeable is the potential equipment of every mind. This assumption in consequence leads to the acceptance of the inner mind system of representation – a certain language of thoughts or ideas common to all minds. If, as Leibniz claims, the mind always thinks, from this it results that it thinks in a certain inner language. This inner mind system of representation corresponds with the whole aperception of the thinking substance – the soul and the mind. This does mean that all perceptions are conscious. The mind perceives the world according to its structure which finds its projection in the system of inner mind representation. It is not conscious of the whole structure and is able to interpret only those fragments which are perceived in the act of aperception.

It is difficult to judge if the above interpretation enables one to recreate Leinbitz’s real views connected in particular with the representational character of unconscious perceptions. However, there are arguments which make this interpretation admissible. The first, is the assumption that lowest on the cognitive scale is the unclear knowledge which is connected with unclear ideas; from this it appears that even an unclear idea is some sort of disposition to think and leads to expression which however is not possible to reach from the level of consciousness. The second, is the Leibnizian idea of a universal algorithm which is labeled as the thread of thought – filum cogitationis.^{10} Finally the third, being supplemental to the previous one, is the belief concerning the existence of a language which was created by Adam, a language which is indeed unknown but whose essence consisted of a natural relationship between names and things.^{11} In contemporary research connected with artificial intelligence appears an assumption that “unconscious processes could well be algoritmic, but at a very complicated level that is monstrusly difficult to disentangle in detail”. (Penrose, 1989, 411) However the algorithm itself must operate on something and this something is presumably indicated by Leibniz as the natural order of ideas.

The assumption of the existence of the natural order of ideas which is the inner mind’s system of representation, a certain language of the mind, is in accord with the rational tradition, and although Leibniz develops his theory of the activity of the mind enriching it with the theory of unconscious perceptions, its general assumptions refer however to the representational theory of the mind laid out by Descartes.^{12} We find the source of this theory in the philosophy of Plato. Plato describes thinking as a silent conversation of the mind with itself. The process of thinking is understood as a discourse, which must assume language.^{13} Wittgenstein, analyzing the views of St. Augustine notices that:

“Augustine describes the learning of human language as if the child came into a strange country and did not understand the language of the country; that is, as if it already had a language, only not this one. Or again: as if the child could already think, only not yet speek. And ’think’ would here mean something like ’talk to itself’”. (Wittgenstein 1958, 15e-16e)

Wittgenstein treats this opinion as absurd although it is fully accepted by present-day representational theory. J. Fodor, a representative of this trend of research, writes that

“one cannot learn a language unless one has a language. In particular, one cannot learn a first language unless one already has a system capable of

10 See Marciszewski 1994, see also C, 351, and C, 153-57.

11 Adamic or paradisaical language was the subject of great speculation in the 17th century. Extensive discussion of this issue may be found in: Aarrsleff, 1982, Losonsky, 1992, and Losonsky, 1993.

12 Chomsky, 1990, 629-30.

13 See: Plato, Theaetetus, 189e-190a. Z. Vendler in his article Wordless Thoughts (Vendler, 1977, 29-30), by interpreting this passage of the dialogue, maintains that Plato treats thinking as a conversation conducted by the mind in some ethnic language, for example Greek. Such interpretation seems to be controversial compared with the fact that, for Plato, language is secondary in relation to thinking itself and constitutes only the initial phase of the learning process. See Plato, Letter VII, see also F. Sontag 1954, 823-830.

representing the predicates in that language and their extensions. And, on pain of circularity, that system cannot be the language that is being learned. But first languages are learned. Hence, at least some cognitive operations are carried out in languages other than natural languages.” (Fodor, 1976, 64)

The assumptions of the existence of the inner language of thought is the starting point for many theories of representation developing at present, among which the theory of Noam Chomsky has received the widest acclaim.

Let us turn now to the sequence of discoveries singled out by Leibniz which is a certain psychological order and which Leibniz calls the history of our discoveries, different with different people. (NE, IV, vii, §9) Reaching aperception or reflection assumes a proper act of dispositions leading to thinking of this or that fragment of reality. Disposition, as we have written above, requires a certain impulse; Leibniz presumably would say: a certain anxiety to such reflective activity. These impulses which usually come from the external senses enable one, when they are properly strengthened, to reach the knowledge of a given thing which may be revealed on many levels of learning. Whichever level of cognitive advancement a given mind has reached, it can be evaluated at the moment this knowledge is externalized. It is unquestionable that language plays here a first rank role. Revealing the content of the mind may be possible through art or an act of construction, a gesture or behavior, but it most fully reveals itself through natural language. One can obviously recall as an argument here the fact that language is the result of aspiration of the thinking substance. It is enough to state here that, by characterizing various levels of learning, Leibniz refers in fact to the language itself. Necessary premises are supplied here by the classification of the learning states set forth in Meditations of Knowledge, Truth, and Ideas (GP, VI, 422-26; L, 291-95) which is as follows:

Knowledge is either obscure or clear. (Est ergo cognito vel obscura vel clara). Clear knowledge (clara) is either distinct (distincta) or confused (confusa). We deal with distinct knowledge when we have at our disposal “means” which enable us to recognize the thing represented (by them). Clear and distinct (clara et distincta) knowledge consists of the ability to enumerate features (notae) sufficient to recognize things. Clear and confused (clara et confusa) knowledge occurs then when we are not able to enumerate one by one those features (notae) which results in us being able to distinguish a given thing from others. Leibniz employs here the example of colors, tastes and smells, which are in fact distinguishable thanks to the evidence of the senses but not through features which are possible to be worked out (non vero notis enuntiabilibus).

Clear and distinct knowledge further divides into adequate and inadequate. Adequate knowledge consists of the ability to enumerate all features (notae) constituting the analyzed thing among which each feature individually is known also distinctly and clearly. If any of these features does not meet the criterion of distinctness then the knowledge is inadequate.

Let us draw our attention to the fact that the criterion of the ability to create definitions becomes essential in distinguishing particular learning states. Most difficult to interpret is the lowest state of knowledge (cognitio obscura) because examples indicated by Leibniz do not explain much here. It seems that obscure knowledge indicates a lack of whatever ability to define. This is a knowledge which we have at our disposal only on the level of unconscious perceptions and as such it is not verbalized. We can interpret this type of knowledge as implicit knowledge, which is the nucleus of other learning states of the mind. We can describe the remaining levels of learning on the basis of the ability to create various kinds of definitions. For example, clear and confused knowledge (cognitio clara et confusa) reveals itself at least as the skill to create an ostensive definition of the object whose idea we possess inside. Partial definitions indicate clear and distinct (cognitio clara et distincta); nominal definitions indicate clear, distinct and inadequate knowledge (cognitio clara, distincta et inadequata); and real definitions indicate clear, distinct and adequate knowledge. We should however stipulate that this assignment illustrates only in a very large approximation the possibility of revealing the particular learning states of a thinking mind. These symptoms of what is invisible and takes place inside of the thinking mind itself are external, possible to observe and measure. Leibniz does not claim that we know only so much as we are able to put into words; this is contradicted by the assumption of the existence of obscure knowledge which is non-verbalizable. However it can be assumed that all remaining forms of knowledge are possible to be reduced to the ability of expressing them in a language; that is, some external language.

Here a question arises: in what relationship to the inner language of thoughts does the natural language in all its variants remain? There is no doubt whatsoever that according to Leibniz “speech expresses thoughts and truths” (GP VII, 263; L, 207). Language is the result which must have its source and for Leibniz this is the permanent and invariable order of ideas. According to him, natural languages have historical character and therefore are carriers of the order of discoveries. This order has a historical as well as psychological dimension. Language is vested only in a human being who is secondary to the mental structures, but consists of one of the most important means of expression of particular learning states of individual users, societies and the whole of Mankind. Language expresses these states thanks to the fact that there exists a certain analogy between all which is enclosed in the language and all that is contained in thoughts. Language refers directly to all which constitutes its source, to the inner world of ideas. Reference to the inner system of representation enables a progressive discovery of new areas which are gradually elevated to the level of consciousness thanks to the capabilities of their verbalization.

The problem of translatability

In this context one is astounded by Leibniz’s strong conviction expressed in Unvorgreiffliche Gedancken that every language, even most limited, is able to express everything. He admittedly notices a certain difficulty, but it is for him rather a matter of form and statistics and not a matter of content because although everything may be expressed through paraphrases or description, the length of the utterance results in the fact that both the speaker and the listener lose the whole pleasure and the proper distribution of accents because it takes up the mind for too long. (UG, §59)

Every language can express everything: this means that it is possible to express all levels of learning accessible to the mind in any language with the assistance of means of expression attainable in that language. But such an approach to language leads to an assumption that it is so to say timeless and is equivalent to the language of the mind. External language as perfect as the natural one reflecting the order of ideas given by God does not exist. Leibniz writes about “any” language having in mind “ethnic” languages. Assuming their timeless character would contradict the thesis of the historical features of natural languages.

The problem of “expressing” is for Leibniz resolved in this context into appropriate rules of language translation while he ignores fully the question of cultural distinction or the level of civilizational development of particular societies. These issues were of course a topic of consideration for 17th century science that was due to the great openness of Europe to the world at that time. Leibniz himself was known for his great fascination with Chinese culture. Therefore, one is surprised at the simplified approach to this so very complicated matter especially when we compare it with Leibniz’s views on the problem of the origin of human speech. Since the original image of the world is imprinted mainly in onomatopoeic words, let us remember that according to Leibniz, primitive peoples had more instinct than reason. (D, IV, ii, 187). How can it then be possible to establish and coordinate the rules of language translation as well as, in a wider sense, the rules of cultural translation? We could ask: is such coordination of the diametrically different states of learning of Man at various levels of civilizational development at all possible?

Leibniz was not ready to give an answer to a question formulated in such a way. He also did not notice a certain contradiction occurring at the theoretical level. The theory of innate ideas assumes the invariability of the equipment of the mind, which means that from the very beginning of his existence, Man had at his disposal a full set of ideas projecting the universe and the changes taking place there. That which determined and still determines the individual character of every mind was the difference in the scale of activated learning dispositions accepted by Leibniz. This difference should also reveal itself in language. Of course this is possible to observe: it includes all verbal behavior of the user of the language, it also enables research of the consecutive stages of his intellectual development. This difference also manifests itself in full language competence through stylistic difference, richness of vocabulary, grammatical competence, etc. In the historical perspective, changes in the domain of the same language take place mainly in the lexical system but also do not exclude grammar. Finally, we have differences between national languages. If, then, an indispensable condition of Leibniz’s principle of individualization is the difference in the scale of learning dispositions of the mind which reveal themselves at the language level – and thanks to it become measurable – how then can one can at the same time assume that what is expressed in the individual language of every user is reciprocally translatable? Since there are no two identical minds, then there are no two identical languages. We could further inquire about the translatability of different language systems, language in its original form and the German language of the 17th and 20th centuries. The only answer is the Leibnizian idea of prior established harmony thanks to which every change taking place in any given substance, and every movement of the mind results in the change of the remainder. But this is not a satisfying answer.

Leibniz, by accepting the fact that languages are the best mirror of the human mind, based this assumption on developed language systems known to him which he researched throughout many years. The conclusions concerned this material which was preserved in written accounts, manuscripts, documents and other sources. He was not able however to interpret facts, as evidenced by his comments concerning languages of the newly discovered tribes: “they differ so much among themselves and are so different from us that we can say that these are a completely different breed”. (AA, I, vii, 399) One may presume that Leibniz had fundamental doubts here and it is not clear if he would admit without hesitation that in these languages also, everything could be expressed. Conclusions educed by contemporary social anthropology rather rule out the possibility of such an approach. For it is accepted that “language in its primitive forms ought to be regarded and studied against the background of human activities and as a mode of human behaviour in practical matters. We have to realize that language originally among primitive, non-civilized peoples was never used as a mere mirror of reflected thought. The manner (...) in which the author of a book, or a papyrus or a hewn iscription has to use it, is a very far-fetched and derivative function of a language. In this, language becomes a condensed piece of reflexion, a record of fact or thought. In its primitive uses language functions as a link in concerted human activity (...). It is a mode of action and not an instrument of reflexion”. (Malinowski, 1956, 312)

The above arguments may lead to the conclusion that there exist essential problems impeding or making the translation of a primitive language to another developed language system impossible. One essential obstacle is, among other things, the reconstruction of the semantic system which is entangled in the pragmatic context.^{14} The problem of translatability is one of the main objections formulated against the contemporary theory of representation. It is pointed out that every description of a learning state of another mind is inadequate due to the lack of objective rules of translation as well as difficulties in establishing the tools of measurement of the accuracy of such translation. Questions of cultural translatability, of the learning states of people with damaged brains, and of children in the pre-verbal phase are also raised.^{15}

However, we may assume that all human language systems remain in a certain relationship with the same ideal system of representation projecting the order of the universe and that this system attempts to decipher Man. For Leibniz assumes that

“all individual created substances, indeed, are different expressions of the same universe and the same universal cause, God. But these expressions vary in perfection as do different representations or perspectives of the same city seen from different points”. (L, 269)

Human languages are the best observable result of the learning activity of the thinking substance. Because they exist there also must exist justification

14 See Malinowski, 1956.

15 See Churchland, Smith Churchland, 1990, 301-302.

for this existence. This constitutes the previously mentioned inner system of expression. The imperfectness of our language fully renders the complexity of the learning processes taking place in our mind. But language refers only to those areas of thinking which are conscious thinking. Leibniz was aware of this fact, and its consequence was his theory of the mind whose essence is the assumption of the gradation of consciousness.^{16} Assuming that the mind always thinks, Leibniz clearly indicated that all which we are unable to examine is unconscious thinking, for it is not verbalized. Thus the process of thinking goes beyond language, but language reflects all that which appears to us in the act of consciousness, and in fact is the only tool enabling us to define the scope of the learning dispositions of a thinking mind.

Leibniz’s Writings

AA – Gottfried Wilhleim Leibniz, Sämtliche Schriften und Briefe, herausgegeben von der Preussischen (jetzt Deutschen) Akademie der Wissenschaften zu Berlin, 6 Reihen, Darmstadt 1923, Leipzig, 1983, Berlin, 1950 – (quoted as volumne, part, page).

C – Opuscules et Fragments Inédits de Leibniz, extraits des manuscrits de la bibliothèque royale de Hanovre par Louis Couturat, Paris, 1903, (repr. Hildesheim, 1961).

D – Gothofredi Guiliemi Leibnitii Opera Omnia, nunc primum colleta... par Ludovicius Dudens, 6 vol. Genève, 1767 – (quoted as volumne, part, page).

GP – Die Philosophischen Schriften von G. W. Leibniz, 7 vol., ed.

C. I. Gerhardt. Halle, 1849-1863 (repr. Hildesheim, 1961). L – Gottfried Wilhelm Leibniz – Philosophical Papers and Letters, ed.

L. E. Loemker, 2nd ed. Dordrecht: Reidel, 1969. NE – G. W. Leibniz, New Essays on Human Understanding, transl. and

ed. P. Remnant and J. Bennett, Cambridge: Cambridge University Press, 1981 – (quoted as book, chapter and paragraph).

UG – G. W. Leibniz, Unvorgreiffliche Gedancken, betreffend die Ausübung und Verbesserung der Teutschen Sprache, in G. G. Leibnitii, Collectanea Etymologica, “illustrationi linguarum, veteris, celticae, germanicae, gallicae, aliarum inservientia”, cum praefatione Johannis Georgii Eccardi, Hannover, 1717.

16 See Dippert, 1994, 177-84.

References

Aarsleff, H. (1982) Leibniz on Locke on Language, in Aarsleff From Locke

to Saussure, Athlone, London. Chomsky, N. (1990) On the Nature, Use and Acquisition of Language, in

Mind and Cognition, Blackwell, Oxford, Cambridge. Churchland, P. M., Smith Churchland, P. (1990) Stalking with the Wild

Epistemic Engine, in Mind and Cognition, Blackwell, Oxford, Cambridge.

Dascal, M. (1987) Leibniz, Language, Signs and Thought, John Benjamins

Publishing Company, Amsterdam//Philadelphia. Dippert, R.R. (1994) Leibniz’s Discussion of Obscure, Confused, and

Inadequate Ideas: Its Relevance for Contemporary Cognitive Science, in Leibniz und Europa, VI Internationaler Leibniz-Kongreß, Hannover.

Fodor, J. A. (1976) The Language of Thouhgt, The Harvester Press Limited. Gadamer, H.G. (1967) Mensch und Sprache, in Gadamer, H. G. “Kleine

Schriften” vol. I, J. C. B. Mohr, Tübingen. Hacking, I. (1975) Why does Language Matter to Philosophy, Cambridge

University Press, Cambridge, London, New York, Melbourne. Jolley, N. (1984) Leibniz and Locke: A Study of the “New Essays on Human

Understanding”, Clarendon Press, Oxford. Jolley, N. (1990) The Light of the Soul, Clarendon Press, Oxford. Losonsky, M. (1992) Leibniz’s Adamic Language of Thought, in “Journal of

the History of Philosophy”, vol. XXX, October 1992, Nb 2. Losonsky, M. (1993) Representationism and Adamicism in Leibniz,

in Leibniz and Adam, ed. Dascal, M., Yakira, E., University Publishing Project Ltd, Tel Aviv.

Malinowski, B. (1931) Culture, in Encyclopaedia of the Social Sciences,

vol. 4, New York, pp. 621-45. Malinowski, B. (1951) The Problem of Meaning in Primitive Languages, in

Ogden, C. K, Richards, I. A., The Meaning of Meaning, Harcourt, Brace and Co. Inc., New York, London.

Marciszewski, W. (1994) Why Leibniz should not have believed in “Filum Cogitationis”, in Leibniz und Europa, VI Internationaler Leibniz-Kongreß, Hannover.

Penrose, R. (1989) The Emperor’s New Mind, Oxford University Press,

Oxford, New York, Melbourne. Sontag, F. (1954) The Platonist Concept of Language, in “The Journal of

Philosophy”, LI//25. Vendler, Z. (1977) Wordless Thoughts, in Language and Thought

Antropological Issues, Mouton Publishers, Hague, Paris. Wittgenstein, L. (1958) Philosophical Investigations, transl. Anscombe,

G. E. M., Basil Blackwell, Oxford.